# Self-embedding similitudes of Bedford-McMullen carpets with dependent ratios

### v2

@@ 1. Introduction
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet with $\frac{\log n}{\log m}\in\mathbb{Q}$ and let $f$ be a similitude sending $K$ into itself. If $K$ is not contained in any line, then $\frac{\log\lambda}{\log n}\in\mathbb{Q}$, where $\lambda$ denotes the contraction ratio of $f$.
@@ 2.1. Bedford-McMullen carpets
Therefore, there must exist some $t_{0}$ such that
@@
The existence of $R$ can be deduced in a similar way we choose $\widetilde{R}$. Let $(a_{1},b_{1})$ be the left endpoint of the original $\ell$ (before the truncation). Let $x^{\prime}_{0}:=\min\{\frac{i}{n^{k}}:\frac{i}{n_{k}}\geq a_{1}\}$ and $y^{\prime}_{0}$ be such that $(x^{\prime}_{0},y^{\prime}_{0})\in\ell$. Then define recursively that
@@
    For every $p,t\geq 1$ and every basic open interval $U$ of $K^{y}_{p}$, $U\cap K^{y}_{p}$ contain $\prod_{k=p+1}^{p+t}\#I_{y_{k}}$ many basic open intervals of $K^{y}_{p+t}$.
@@
$K^{y}_{p+t}\cap U=\bigcup\Big{\{}\Big{(}a+\sum_{k=p+1}^{p+t}\frac{x_{k}}{n^{k}},\Big{(}a+\sum_{k=p+1}^{p+t}\frac{x_{k}}{n^{k}}\Big{)}+n^{-p-t}\Big{)}:x_{k}\in I_{y_{k}}\Big{\}},$
@@ 2.2. Deleted-digit sets
Let $n\geq 2$ be an integer. For any non-empty $\mathcal{D}\subset\{0,1,\ldots,n-1\}$, we write $E(n,\mathcal{D})$ to be the self-similar set associated with the IFS $\{\frac{x+i}{n}:i\in\mathcal{D}\}$ on $\mathbb{R}$. In other words,
@@
By a renowned theorem of Schief [17], $E(n,\mathcal{D})$ has non-empty interior if and only if $\operatorname{dim_{H}}E(n,\mathcal{D})=1$. In particular, if $\operatorname{dim_{H}}E(n,\mathcal{D})<1$ then it must be totally disconnected. It is easy to check by induction that $E_{k}(n,\mathcal{D})$ is a union of intervals of the form $[\frac{i}{n^{k}},\frac{i+1}{n^{k}}]$, and each pair of such intervals are either adjacent (that is, they share a common endpoint) or disjoint.
@@
Without loss of generality, assume that $h$ is orientation-preserving. If $y\in\pi_{2}(K)$ has more than one expansion with coefficients in $J$, $y\in\bigcup_{k\geq 1}\frac{\mathbb{Z}}{m^{k}}\cap[0,1]$. In this case, it is easy to see that $K^{y}$ is a finite union of scaled copies of $K^{0}$ and $K^{1}$, say $K^{y}=\bigcup_{s\in\Omega}K^{0,s}\cup\bigcup_{t\in\Gamma}K^{1,t}$, where $\Omega,\Gamma\subset\mathbb{Z}$ and
@@
Case 1: $h^{s}(B)$ intersects exactly one connected component of $K^{y}_{p_{0}+t}$, say $B^{\prime}$. In this case, $h^{s}(B\cap K^{y})\subset B^{\prime}\cap K^{y}$. Since $B$ consists of $M$ basic open intervals of $K^{y}_{p_{0}}$ and their intersections with $K^{y}$ are distinct from one another by only a translation (recall Lemma 2.3(1)), $h^{s}(B\cap K^{y})$ contains a $\rho^{s}n^{-p_{0}}$-separated subset of cardinality $M$. Since $\rho^{s}n^{-p_{0}}>n^{-p_{0}-t}$, every basic open interval of $K^{y}_{p_{0}+t}$ in $B^{\prime}$, which has length $n^{-p_{0}-t}$, contains at most one point in that subset. Thus by the maximality of $M$, $B^{\prime}$ contains exactly $M$ basic open intervals of $K^{y}_{p_{0}+t}$. Recall that $0<\mathcal{H}^{\alpha}(K^{y})<\infty$. Writing $c$ to be the $\alpha$-Hausdorff measure of the intersection of $K^{y}$ and any basic open interval in $K^{y}_{p_{0}+t}$, we see from Lemma 2.3(1) that $0<c<\infty$ and $\mathcal{H}^{\alpha}(B^{\prime}\cap K^{y})=Mc$. On the other hand, by Lemma 2.3(2), each basic open interval in $B$ contains exactly $\prod_{k=p_{0}+1}^{p_{0}+t}\#I_{y_{k}}=N^{t}$ ones in $K^{y}_{p_{0}+t}$. So $\mathcal{H}^{\alpha}(B\cap K^{y})=MN^{t}c$ and
@@
Case 2: $h^{s}(B)$ intersects at least two connected components of $K^{y}_{p_{0}+t}$. In this case, $h^{s}(B)$ contains a hole between those components. More precisely, there exists some integer $i$ such that $(\frac{i}{n^{p_{0}+t}},\frac{i+1}{n^{p_{0}+t}})\cap K^{y}=\varnothing$ and $(\frac{i}{n^{p_{0}+t}},\frac{i+1}{n^{p_{0}+t}})\subset h^{s}(B)$. Write $V:=(\frac{i}{n^{p_{0}+t}},\frac{i+1}{n^{p_{0}+t}})$. Comparing the length, we can find either exactly one or two adjacent basic open intervals of $K^{y}_{p_{0}}$ in $B$ of which the images under $h^{s}$ meet $V$. If $V\subset h^{s}(\widetilde{I})$ for only one basic open interval $\widetilde{I}$ in $B$ (see Figure 1(A)), then, since $h^{s}(\widetilde{I}\cap K^{y})\subset K^{y}$, $h^{s}(\widetilde{I}\cap K^{y})\cap V=\varnothing$. This in turn tells us that $\widetilde{I}$ contains a gap of length $\geq|h^{-s}(V)|$. Therefore,
@@
Now, suppose there are two basic open intervals $\widetilde{I}_{1},\widetilde{I}_{2}$ of $K^{y}_{p_{0}}$ in $B$ such that $V\cap h^{s}(\widetilde{I}_{i})\neq\varnothing$, $1\leq i\leq 2$. Without loss of generality, assume that $\widetilde{I}_{1}$ is to the left of $\widetilde{I}_{2}$. Denote by $a_{i}$ the left endpoint of $h^{s}(\widetilde{I}_{i})$. See Figure 1(B) for an illustration.
@@
Let $K,N,y,h$ be as in Proposition 2.6. Then there are left and right isolated points $a,b$ of $K^{y}$ such that $h(a),h(b)$ (or $h(b),h(a)$ if $h$ is orientation-reversing) are also left and right isolated points of $K^{y}$. Moreover, we can guarantee that $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$ for all large $t$, where $R_{t},R^{\prime}_{t}$ are two level-$t$ rectangles with $(a,y)\in\varphi_{R_{t}}(K)$ and $(b,y)\in\varphi_{R^{\prime}_{t}}(K)$, respectively.
@@
Similarly as before, let $\rho$ be the contraction ratio of $h$. By Proposition 2.6, $\rho=n^{-p/q}$ for some positive integers $p,q$. Applying the argument in the proof of Proposition 2.6 to $s=q$, $t=p$ and $\varepsilon=0$, we see that Case 2 in that proof is impossible because: (1) the right hand side of (2.11) now equals $1$, i.e., $\widetilde{I}$ contains a gap of the same length as itself, which contradicts that $\widetilde{I}\cap K^{y}\neq\varnothing$; (2) the right hand side of (2.12) now equals 00, i.e., $|h^{q}(\widetilde{I_{1}}\cap K^{y})|=0$, which contradicts that $\mathcal{H}^{\alpha}(h^{q}(\widetilde{I_{1}}))>0$. On the other hand, for Case 1 in that proof, we now have $\mathcal{H}^{\alpha}(h^{q}(B\cap K^{y}))=\mathcal{H}^{\alpha}(B^{\prime}\cap K^{y})$. We claim that
@@
Otherwise, since $h^{q}(B\cap K^{y})\subset B^{\prime}\cap K^{y}$, there exists some $x\in B^{\prime}\cap K^{y}\setminus h^{q}(B\cap K^{y})$. So we can find a large $t$ and a basic open interval $W$ of $K^{y}_{t}$ such that $x\in W$ but $W\cap h^{q}(B\cap K^{y})=\varnothing$. However, since $0<\mathcal{H}^{\alpha}(K^{y})<\infty$, we have again by Lemma 2.3(1) that $\mathcal{H}^{\alpha}(W\cap K^{y})>0$. Thus
@@
For the second one, since $y$ has a unique expansion, $R_{t}$ and $R^{\prime}_{t}$ are unique for all $t$. Moreover, $y$ belongs to the interior of both of $\pi_{2}(R_{t})$ and $\pi_{2}(R^{\prime}_{t})$, which implies that $\pi_{2}(R_{t})=\pi_{2}(R^{\prime}_{t})$.
Now, suppose $y$ has more than one expansion. Similarly as in the proof of Proposition 2.6, we can write $K^{y}=\bigcup_{s\in\Omega}K^{0,s}\cup\bigcup_{t\in\Gamma}K^{1,t}$, where $\Omega,\Gamma\subset\mathbb{Z}$ and
$K^{0,s}:=\frac{K^{0}+s}{n^{k}}\quad\text{and}\quad K^{1,s}:=\frac{K^{1}+t}{n^{k}}$
for some large integer $k$. By Baire’s theorem again, we may assume without loss of generality that some $K^{0,s}$ contains an interior part of $h(K^{y})$. In particular, there is a scaled copy of $K^{0,s}$, say $E=\frac{K^{0}+z}{n^{p}}\subset K^{0,s}$, such that $h(E)\subset K^{0,s}$. Note that
$h(E)\subset K^{0,s}\Longleftrightarrow h\Big{(}\frac{K^{0}+z}{n^{p}}\Big{)}\subset\frac{K^{0}+s}{n^{k}}.$
We can easily reformulate this as $g(K^{0})\subset K^{0}$ for some similitude $g$ with contraction ratio $\rho n^{k-p}$. Since 00 has a unique expansion, as in the first part of this proof, there are nondegenerate open intervals $U,V$ such that $g(U\cap K^{0})=V\cap K^{0}\neq\varnothing$ (recall (2.13)). In particular, $g$ maps all the one-sided isolated points of $K^{0}$ in $U$ to those of $K^{0}$ in $V$. Since the inclusion $g(K^{0})\subset K^{0}$ is just a reformulation of $h(E)\subset K^{0,s}$, $h$ maps all one-sided isolated points of $K^{0,s}$ in $U^{\prime}$ (for some $U^{\prime}$) to one-sided isolated points of $K^{0,s}$ in $V^{\prime}$ (for some $V^{\prime}$). It remains to prove that these points are also one-sided isolated points of $K^{y}$.
@@
(2.14) $K^{0,s}\cap V^{\prime}=K^{y}\cap V^{\prime}.$
In fact, if $t\neq s$ for all $t\in\Gamma$, then we already have (2.14). If some $t$ equals $s$, then $K^{0,s}$ and $K^{1,t}=K^{1,s}$ may overlap. However, note that $\operatorname{dim_{H}}K^{0,s}\geq\operatorname{dim_{H}}K^{y}\geq\operatorname{dim_{H}}K^{1,s}$ (recall that $h(K^{y})\subset K^{0,s}$), we have either $K^{1,s}=K^{0,s}$ or these two deleted-digit sets have different patterns (in particular, $K^{0,s}\neq K^{1,s}$). So there is a small open interval $V^{\prime\prime}\subset V^{\prime}$ such that $(K^{0,s}\cap V^{\prime\prime})\cap K^{1,s}=\varnothing$ but $K^{0,s}\cap V^{\prime\prime}\neq\varnothing$. Replacing $V^{\prime}$ with $V^{\prime\prime}$ gives us (2.14). In particular, all left (resp. right ) isolated points of $K^{0,s}$ in $V^{\prime}$ are left (resp. right) isolated points of $K^{y}$.
Next, if a one-sided isolated point $a$ of $K^{0,s}$ in $U^{\prime}$ is not a one-sided isolated point of $K^{y}$, there is a sequence $\{x_{n}\}\subset K^{y}$ such that $x_{n}\to a$ from the left or from the right. Then $h(x_{n})\to h(a)$ from one direction. Since $h(x_{n})\in K^{y}$, this contradicts the fact that $h(a)$ is a one-sided isolated point of $K^{y}$ in $V^{\prime}$, which completes the proof. ∎

### v3

@@ 1. Introduction
Let $K=K(n,m,\Lambda)$ be a Bedford-McMullen carpet with $\frac{\log n}{\log m}\in\mathbb{Q}$ and let $f$ be a similitude sending $K$ into itself. If $K$ is not contained in any line, then $\frac{\log\lambda}{\log n}\in\mathbb{Q}$, where $\lambda$ denotes the similarity ratio of $f$.
@@ 2.1. Bedford-McMullen carpets
Therefore, there must exist some $10\leq t_{0}\leq n^{p+2}-$ 1 such that
@@
The existence of $R$ can be deduced in a similar way we choose $\widetilde{R}$. Let $(a_{1},b_{1})$ be the left endpoint of the original $\ell$ (before the truncation). Pick $k$ large enough such that $n^{-k}$ is much smaller than $|\pi_{1}(\ell)|$. Let $x^{\prime}_{0}:=\min\{\frac{i}{n^{k}}:\frac{i}{n_{k}}\geq a_{1}\}$ and $y^{\prime}_{0}$ be such that $(x^{\prime}_{0},y^{\prime}_{0})\in\ell$. Then define recursively that
@@
    For every $p,t\geq 1$ and every basic open interval $U$ of $K^{y}_{p}$, $U\cap K^{y}_{p}$ contains $\prod_{k=p+1}^{p+t}\#I_{y_{k}}$ many basic open intervals of $K^{y}_{p+t}$.
@@
$K^{y}_{p+t}\cap U=\bigcup\Big{\{}\Big{[}a+\sum_{k=p+1}^{p+t}\frac{x_{k}}{n^{k}},\Big{(}a+\sum_{k=p+1}^{p+t}\frac{x_{k}}{n^{k}}\Big{)}+n^{-p-t}\Big{]}:x_{k}\in I_{y_{k}}\Big{\}}\setminus\{a,a+n^{-p}\},$
@@ 2.2. Deleted-digit sets
Let $n\geq 2$ be an integer. For any non-empty set $\mathcal{D}\subset\{0,1,\ldots,n-1\}$, we write $E(n,\mathcal{D})$ to be the self-similar set associated with the IFS $\{\frac{x+i}{n}:i\in\mathcal{D}\}$ on $\mathbb{R}$. In other words,
@@
It is clear that $E(n,\mathcal{D})$ has non-empty interior if and only if #$\#\mathcal{D}=n$. In particular, if $\operatorname{dim_{H}}E(n,\mathcal{D})<1$ then it must be totally disconnected. It is easy to check by induction that $E_{k}(n,\mathcal{D})$ is a union of intervals of the form $[\frac{i}{n^{k}},\frac{i+1}{n^{k}}]$, and each pair of such intervals are either adjacent (that is, they share a common endpoint) or disjoint.
@@
Considering $h^{2}$ if necessary, we may assume that $h$ is orientation-preserving. If $y\in\pi_{2}(K)$ has more than one expansion with coefficients in $J$, $y\in\bigcup_{k\geq 1}\frac{\mathbb{Z}}{m^{k}}\cap[0,1]$. In this case, it is easy to see that $K^{y}$ is a finite union of scaled copies of $K^{0}$ and $K^{1}$, say $K^{y}=\bigcup_{s\in\Omega}K^{0,s}\cup\bigcup_{t\in\Gamma}K^{1,t}$, where $\Omega,\Gamma\subset\mathbb{Z}$ and
@@
Case 1: $h^{s}(B)$ intersects exactly one connected component of $K^{y}_{p_{0}+t}$, say $B^{\prime}$. In this case, $h^{s}(B\cap K^{y})\subset B^{\prime}\cap K^{y}$. Since $B$ consists of $M$ basic open intervals of $K^{y}_{p_{0}}$ and their intersections with $K^{y}$ are distinct from one another by only a translation (recall Lemma 2.3(1)), one can find a $\rho^{s}n^{-p_{0}}$-separated subset of $h^{s}(B\cap K^{y})$ that has cardinality $M$. Since $\rho^{s}n^{-p_{0}}>n^{-p_{0}-t}$, every basic open interval of $K^{y}_{p_{0}+t}$ in $B^{\prime}$, which has length $n^{-p_{0}-t}$, contains at most one point in that subset. Thus by the maximality of $M$, $B^{\prime}$ contains exactly $M$ basic open intervals of $K^{y}_{p_{0}+t}$. Recall that $0<\mathcal{H}^{\alpha}(K^{y})<\infty$. Writing $c$ to be the $\alpha$-Hausdorff measure of the intersection of $K^{y}$ and any basic open interval in $K^{y}_{p_{0}+t}$, we see from Lemma 2.3(1) that $0<c<\infty$ and $\mathcal{H}^{\alpha}(B^{\prime}\cap K^{y})=Mc$. On the other hand, by Lemma 2.3(2), each basic open interval in $B$ contains exactly $\prod_{k=p_{0}+1}^{p_{0}+t}\#I_{y_{k}}=N^{t}$ ones in $K^{y}_{p_{0}+t}$. So $\mathcal{H}^{\alpha}(B\cap K^{y})=MN^{t}c$ and
@@
Case 2: $h^{s}(B)$ intersects at least two connected components of $K^{y}_{p_{0}+t}$. In this case, $h^{s}(B)$ contains a hole between those components. More precisely, there exists some integer $i$ such that $(\frac{i}{n^{p_{0}+t}},\frac{i+1}{n^{p_{0}+t}})\cap K^{y}=\varnothing$ and $(\frac{i}{n^{p_{0}+t}},\frac{i+1}{n^{p_{0}+t}})\subset h^{s}(B)$. Write $V:=(\frac{i}{n^{p_{0}+t}},\frac{i+1}{n^{p_{0}+t}})$. Comparing the length, we can find either exactly one or two adjacent basic open intervals of $K^{y}_{p_{0}}$ in $B$ of which the images under $h^{s}$ meet $V$.
Case 2.1: $V\subset h^{s}(\widetilde{I})$ for only one basic open interval $\widetilde{I}$ in $B$ (see Figure 1(A)). Then, since $h^{s}(\widetilde{I}\cap K^{y})\subset K^{y}$, $h^{s}(\widetilde{I}\cap K^{y})\cap V=\varnothing$. This in turn tells us that $\widetilde{I}$ contains a gap of length $\geq|h^{-s}(V)|$. Therefore,
@@
Case 2.2: There are two basic open intervals $\widetilde{I}_{1},\widetilde{I}_{2}$ of $K^{y}_{p_{0}}$ in $B$ such that $V\cap h^{s}(\widetilde{I}_{i})\neq\varnothing$, $1\leq i\leq 2$. Without loss of generality, assume that $\widetilde{I}_{1}$ is to the left of $\widetilde{I}_{2}$. Denote by $a_{i}$ the left endpoint of $h^{s}(\widetilde{I}_{i})$. See Figure 1(B) for an illustration.
@@
Let $K,N,y,h$ be as in Proposition 2.6. Then there are left and right isolated points $a,b$ of $K^{y}$ such that $h(a),h(b)$ (or $h(b),h(a)$ if $h$ is orientation-reversing) are also left and right isolated points of $K^{y}$. Moreover, we can guarantee that $\pi_{2}(R_{k})=\pi_{2}(R^{\prime}_{k})$ for all large $k$, where $R_{k},R^{\prime}_{k}$ are two level-$k$ rectangles with $(a,y)\in\varphi_{R_{k}}(K)$ and $(b,y)\in\varphi_{R^{\prime}_{k}}(K)$, respectively.
@@
Similarly as before, let $\rho$ be the contraction ratio of $h$. By Proposition 2.6, $\rho=n^{-p/q}$ for some positive integers $p,q$. Applying the argument in the proof of Proposition 2.6 to $s=q$, $t=p$ (so $\varepsilon=s\cdot\frac{\log\rho}{\log n}-t=0$), we see that Case 2 in that proof is impossible because: (1) in Case 2.1, the right hand side of (2.11) now equals $1$, i.e., $\widetilde{I}$ contains a gap of the same length as itself, which contradicts that $\widetilde{I}\cap K^{y}\neq\varnothing$; (2) in Case 2.2 the right hand side of (2.12) now equals 00, i.e., $|h^{q}(\widetilde{I_{1}}\cap K^{y})|=0$, which contradicts that $\mathcal{H}^{\alpha}(h^{q}(\widetilde{I_{1}}\cap K^{y}))>0$. On the other hand, for Case 1 in that proof, we now have $\mathcal{H}^{\alpha}(h^{q}(B\cap K^{y}))=\mathcal{H}^{\alpha}(B^{\prime}\cap K^{y})$. We claim that
@@
Otherwise, since $h^{q}(B\cap K^{y})\subset B^{\prime}\cap K^{y}$, there exists some $x\in B^{\prime}\cap K^{y}\setminus h^{q}(B\cap K^{y})$. So we can find a large $t$ and a basic open interval $W$ of $K^{y}_{t}$ such that $x\in W\cap B^{\prime}$ but $W\cap h^{q}(B\cap K^{y})=\varnothing$. However, since $0<\mathcal{H}^{\alpha}(K^{y})<\infty$, we have again by Lemma 2.3(1) that $\mathcal{H}^{\alpha}(W\cap K^{y})>0$. Thus
@@
For the second one, since $y$ has a unique expansion, $R_{k}$ and $R^{\prime}_{k}$ are unique for all $k$. Moreover, $y$ belongs to the interior of both of $\pi_{2}(R_{k})$ and $\pi_{2}(R^{\prime}_{k})$, which implies that $\pi_{2}(R_{k})=\pi_{2}(R^{\prime}_{k})$.
Now, suppose $y$ has more than one expansion. Similarly as in the proof of Proposition 2.6, we can write $K^{y}=\bigcup_{i\in\Omega}K^{0,i}\cup\bigcup_{j\in\Gamma}K^{1,j}$, where $\Omega,\Gamma\subset\mathbb{Z}$ and
$K^{0,i}:=\frac{K^{0}+i}{n^{k}}\quad\text{and}\quad K^{1,j}:=\frac{K^{1}+j}{n^{k}}$
for some large integer $k$. By Baire’s theorem again, we may assume without loss of generality that some $K^{0,i}$ contains an interior part of $h(K^{y})$. In particular, there is a scaled copy of $K^{0,i}$, say $E=\frac{K^{0}+z}{n^{c}}\subset K^{0,i}$, such that $h(E)\subset K^{0,i}$. Note that
$h(E)\subset K^{0,i}\Longleftrightarrow h\Big{(}\frac{K^{0}+z}{n^{c}}\Big{)}\subset\frac{K^{0}+i}{n^{k}}.$
We can easily reformulate this as $g(K^{0})\subset K^{0}$ for some similitude $g$ with contraction ratio $\rho n^{k-c}$. Since 00 has a unique expansion, as in the first part of this proof, there are nondegenerate open intervals $U,V$ such that $g(U\cap K^{0})=V\cap K^{0}\neq\varnothing$ (recall (2.13)). In particular, $g$ maps all the one-sided isolated points of $K^{0}$ in $U$ to those of $K^{0}$ in $V$. Since the inclusion $g(K^{0})\subset K^{0}$ is just a reformulation of $h(E)\subset K^{0,i}$, $h$ maps all one-sided isolated points of $K^{0,i}$ in $U^{\prime}$ (for some $U^{\prime}$) to one-sided isolated points of $K^{0,i}$ in $V^{\prime}$ (for some $V^{\prime}$). It remains to prove that these points are also one-sided isolated points of $K^{y}$ (for all large $k$, one can pick $R_{k},R^{\prime}_{k}$ to be those that lie above adjacent to ×$\mathbb{R}\times\{y\}$ and thus $\pi_{2}(R_{k})=\pi_{2}(R^{\prime}_{k})$).
@@
(2.14) $K^{0,i}\cap V^{\prime}=K^{y}\cap V^{\prime}.$
In fact, if $j\neq i$ for all $j\in\Gamma$, then we already have (2.14). If some $j$ equals $i$, then $K^{0,i}$ and $K^{1,j}=K^{1,i}$ may overlap. However, note that $\operatorname{dim_{H}}K^{0,i}\geq\operatorname{dim_{H}}K^{y}\geq\operatorname{dim_{H}}K^{1,i}$ (recall that $h(K^{y})\subset K^{0,i}$), we have either $K^{1,i}=K^{0,i}$ or these two deleted-digit sets have different patterns (in particular, $K^{0,i}\neq K^{1,i}$). So there is a small open interval $V^{\prime\prime}\subset V^{\prime}$ such that $(K^{0,i}\cap V^{\prime\prime})\cap K^{1,i}=\varnothing$ but $K^{0,i}\cap V^{\prime\prime}\neq\varnothing$. Replacing $V^{\prime}$ with $V^{\prime\prime}$ gives us (2.14). In particular, all left (resp. right ) isolated points of $K^{0,i}$ in $V^{\prime}$ are left (resp. right) isolated points of $K^{y}$.
Next, if a one-sided isolated point $a$ of $K^{0,i}$ in $U^{\prime}$ is not a one-sided isolated point of $K^{y}$, there is a sequence $\{x_{n}\}\subset K^{y}$ such that $x_{n}\to a$ from the left or from the right. Then $h(x_{n})\to h(a)$ from one direction. Since $h(x_{n})\in K^{y}$, this contradicts the fact that $h(a)$ is a one-sided isolated point of $K^{y}$ in $V^{\prime}$, which completes the proof. ∎